OpenStudy (anonymous):

Find the equation of the line that passes through (3,4) and m = -4.

OpenStudy (anonymous):

use point slope formula y-y1=m(x-x1) plug and chug

OpenStudy (anonymous):

like the last one: y - 4 = -4(x - 3) and solve for y